OpenStudy (anonymous):

Can u help me with this one?: Write an equation that shows all of the values of theta in radians such that sin (theta)=0.

OpenStudy (mathmale):

Hint: the sine function is periodic; that is, it repeats itself every 2pi radians. Hint #2: sin 0 = 0, sin pi = 0; sin 2pi = 0, and so on. Go ahead and write the expression that shows all of the values of theta in radians such that sin (theta) = 0.

OpenStudy (jdoe0001):

look at you Unit Circle, what angle has sine = 0 ?

OpenStudy (anonymous):

180 degrees and 360 degrees

OpenStudy (anonymous):

I have no clue how to make an equation out of that number

OpenStudy (mathmale):

Yes, similar to sin 0 = 0, sin pi = 0; sin 2pi = 0, and so on. You now have all the info you need to write an expression for the solutions of sin theta = 0. The solution set is 0 plus or minus n*Pi, or 0 plus or minus n*180 deg. Try it. Let n = 0, 1, 2, 3, and so on. Do these angle values make sin (theta) = 0? Note that n can also be negative: n = -1, -2, -3, and so on.

OpenStudy (anonymous):

0

OpenStudy (jdoe0001):

yeap and if you keep on going around pass the 360 degrees say 540, 720 and so on, those will be co-terminal angles to the ones you mentioned and thus will have the same sine = 0 and those are the angles whose sine is 0 :)

OpenStudy (anonymous):

so whats the equation then?

OpenStudy (anonymous):

ugh annoyances of math why do u have to be sooo dang difficult.

OpenStudy (mathmale):

this is probably a new situation for you. Simply write \[\theta \pm n(180 \deg), n={1, 2, 3, 4, ...}\]

OpenStudy (mathmale):

The solution set is 0 plus or minus n*Pi, or 0 plus or minus n*180 deg. this is given symbolically in the equation I typed, immediately above.

OpenStudy (anonymous):

Thank you sooo much....is it possible to help me with another?

OpenStudy (mathmale):

Please post it as a new problem. But first, please apply whatever you can of what you learne3d here to solving this new problem. Thanks.

OpenStudy (jdoe0001):

I'd think something like -> \(\bf sin(\theta)=0;\quad \theta = \pi + n; n \in \mathbb{Z}\)